OpenStudy (anonymous):

HELP PLEASEE! If f(x) = 5x2 – 8x + 13 and g(x) = 6x – 5, what is f(x) + g(x)? 5x2 – 14x + 18 5x2 – 2x + 18 5x2 – 14x + 8 5x2 – 2x + 8

OpenStudy (anonymous):

12

OpenStudy (anonymous):

f(x) = 5x2 – 8x + 13 and g(x) = 6x – 5 f(x) + g(x) = 5x2 – 8x + 13 + 6x – 5 = 5x2 -8x+6x +13-5 do the last step

OpenStudy (anonymous):

I think so.. So would it be 5x^2-2x+8?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Thank you!